A coffee distributor needs to mix a(n) Rift Valley coffee blend that normally sells for $8.60 per pound with a Kona coffee blend that normally sells for $13.20 per pound to create 10 pounds of a coffee that can sell for $12.74 per pound. How many pounds of each kind of coffee should they mix?
Answer by addingup(3677)

I'll call the quantity of RV x. Then the quantity of Kona Blend will be 10-x
:
8.60x+13.20(10-x) = 12.74(10)
8.60x+132-13.20x = 127.40
-4.60 = -4.60
x = 1 You need 1 pound of 8.60 and 10-1= 9 pounds of 13.20
------------------------------
Check:
1 lb at 8.60 = 8.60
and 9 lbs at 13.20 = 118.80
8.60+118.80 = 127.40 and 127.40/10 = 12.74 per pound. Correct.
;
John